Find the probability that sum of the numbers showing on the two dice is 8, given that atleast one
dice doesnt show five.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
of the 36possible outcomesfor rolling two dice, the only one not to be considered here is (5,5). For all other outcomes, at least one die doesn't show a 5.
Also for sum=8, (6,2)(2,6)(3,5)(5,3)(4,4)
p(sum=8/atleast one dont show 5)=5/35=1/7
Thanks and Regards
Shaik Aasif
askIITians faculty